# Capital Asset Pricing Model with Size Factor and Normalizing by Volatility Index

### v1

@@ 6. Conclusion and Further Research
We combined main results of our previous articles [11, 17]. In [11], the main model (CAPM plus linear dependence of $\alpha$ and $\beta$ upon relative size) was used to capture the property that small stocks have, on average, higher risks and returns. In this article, we add to this model the normalization (division by the Volatility Index). The resulting multivariate Markov time series model fits better using real-world market data: Its residuals are better descrbied as IID Gaussian. We state and prove a simple sufficient condition (Theorem 1) for ergodicity. And we make some connections with Stochastic Portfolio Theory, where several other models were proposed trying to capture this size effect.
@@
Appendix
We stated and proved in [11] results on stationarity and ergodicity of a continuous-time analogue of our model, but with constant volatility. We did not state and prove stability result for discrete time. Here, we fill this gap. In the notation of this article, we can assume $V(t)=1$ is constant. This corresponds to $W(t)\equiv 0$. This violates Assumption 1 (the part with positive density). But this does not affect the stationarity proof.
Assumption 4.
$Z(t)\sim\mathcal{N}(0,\sigma^{2})$ IID, and $V(t)=1$.
Under Assumption 4, let us check Assumption (9). We have: $R_{0}(t)=g+Z(t)$. Then we can rewrite the left-hand side of Assumption 23 as
(35) $\begin{split}\mathbb{E}&\left[\ln|\xi|\right]<0,\\ \xi&:=1+a+bR_{0}(t)\sim\mathcal{N}(\mu,\rho^{2}),\\ \mu&:=1+a+bg,\quad\rho:=|b|\sigma.\end{split}$
is shown in Figure 1. Python code is available on GitHub repository asarantsev/normal-log.

### v3

@@ 6. Conclusion and Further Research
We combined main results of our previous articles [13, 22]. In [13], the main model (CAPM plus linear dependence of $\alpha$ and $\beta$ upon relative size) was used to capture the property that small stocks have, on average, higher risks and returns. In this article, we add to this model the normalization (division by the Volatility Index). The resulting multivariate Markov time series model fits better using real-world market data: Its residuals are better descrbied as IID Gaussian. We state and prove a simple sufficient condition (Theorem 1) for ergodicity. And we make some connections with Stochastic Portfolio Theory, adding our model to the collection of proposed models capturing this size effect. We state and prove rigorous results on the capital distribution curve.
We fill two lacunas in our previous research from [13].
First lacuna: Long-term stability results apply to the case of constant volatility from [13], see subsection 4.4 of the current article.
Second lacuna: The curve is order statistics of the normal sample (but with random mean and standard deviation), see subsection 5.3 of the current article, continued below in the Appendix. Such curve reproduces the real-world shape of the captial distribution curve: linear at the upper left end, and concave at the lower right end.
@@
Appendix: Standard Normal Market Curve
To continue subsection 5.3, it remains to study the behavior of the curve (41) using the classic Extreme Value Theory. Most of the definitions and results of this subsection are well-known. We do not even attempt to provide an exhaustive list of references. Instead, we mention classic monograph [23] and a classic textbook [2]. For Poisson point processes on the real line, see another classic monograph [15].
Pick a function $\lambda:\mathbb{R}\to[0,\infty)$ which is locally integrable: $\Lambda(I):=\int_{I}\lambda(x)\,\mathrm{d}x<\infty$ for any bounded interval $I\subseteq\mathbb{R}$.
Definition 5.
A Poisson point process on $\mathbb{R}_{+}$ with intensity or rate $\lambda$ is defined as a random countable subset $\mathcal{N}\subseteq\mathbb{R}$ such that $\mathcal{N}\cap I\sim\mathrm{Poi}(\Lambda(I))$. That is, the random number of points on any bounded interval $I$ is Poisson with mean $\int_{I}\lambda(u)\,\mathrm{d}u$.
In particular, if $\lambda(t)=e^{-t}$, then we can rank points from the rightmost to the left. In other words, the $k$th rightmost point $\mathcal{N}^{+}_{k}$ is well-defined. This point exists, since $\lambda$ is integrable on $[u,\infty)$ for any $u\in\mathbb{R}$. Denote by $\tau_{k}$ the $k$th jump time of the standard Poisson process on $[0,\infty)$: $\tau_{k}-\tau_{k-1}$ are IID exponential with mean 1, where by convention $\tau_{0}:=0$. For fixed $m$, we can also express
(42) $\mathcal{N}^{+}_{k}=-\ln(\tau_{k}),\quad k=1,\ldots,m.$
Similarly, if $\lambda(t)=e^{t}$, then we can rank points from the leftmost to the right. And the formula (42) becomes the formula for the $k$th leftmost point:
(43) $\mathcal{N}^{-}_{k}=\ln\tau_{k},\quad k=1,\ldots,m.$
See also [2, Chapter 8, Exercises 6, 7].
Definition 6.
The (standard) Gumbel distribution $G$ is defined by its cumulative distribution function $\exp(-\exp(-x))$.
It is known, see classic references [2, Theorem 8.3.1, Example 8.3.4] or [23, Chapter 1], that the normal distribution belongs to the Gumbel domain of attraction: Consider the maximum of $n$ IID normal variables:
$M_{n}=\max(X_{1},\ldots,X_{n}),\quad X_{1},X_{2},\ldots\sim\mathcal{N}(0,1)\ \mbox{IID.}$.
After scaling, this maximum converges weakly to Gumbel distribution as $n\to\infty$:
[…]
where $a_{n}>0$ and $b_{n}$ are suitable constants. A common suggestion is
$a_{n}=\frac{1}{\sqrt{2\ln n}},\quad b_{n}=\sqrt{2\ln n}-\frac{\ln(4\pi\ln n)}{2\sqrt{2\ln n}}.$
But the convergence rate for this choice of constants is not the best, as shown in [2, Example 8.3.4]. There are ways to improve this, for example [14]:
$n\varphi(b_{n})=b_{n},\quad a_{n}=1/b_{n},\quad\varphi(u):=\frac{1}{\sqrt{2\pi}}\exp\left[-\frac{u^{2}}{2}\right].$
For any such sequences $(a_{n})$ and $(b_{n})$, we have convergence of top $k$ ranked standardized variables $X_{(1)}>\ldots>X_{(k)}$ to the rightmost $k$ points of $\mathcal{N}$, as $N\to\infty$:
[…]
Similarly and symmetrically, as $N\to\infty$,
[…]
This convergence (45) or (46) follows from [2, Theorem 8.4.2], or [18], or [23, Chapter 4]; see also [2, Chapter 8, Exercises 6–8] and compare with (42) or (43). Recall (42) and plot this Poisson point process with $x$-axis log scale. This represents (up to shift by $b_{N}$ and rescaling by $a_{N}$) the left upper end of the capital distribution curve:
(47) $(\ln k,\mathcal{N}^{+}_{k})=(\ln k,-\ln(\tau_{k})),\,k=1,2,\ldots,m.$
Similarly, recall (43) and plot this Poisson point process with $x$-axis log scale. This represents (up to shift by $-b_{N}$ and rescaling by $a_{N}$) the right lower end of the capital distribution curve:
(48) $(\ln(N+1-k),\mathcal{N}^{-}_{k})=(\ln(N+1-k),\ln(\tau_{k})),\,k=1,2,\ldots,m.$
Lemma 3.
For each $k$, the distribution of $\ln\tau_{k+1}-\ln\tau_{k}$ is exponential with mean $1/k$.
Proof.
From definitions of $\tau_{k}$, we have: $\tau_{k}$ and $\tau_{k+1}-\tau_{k}$ are independent; $\tau_{k}$ has Gamma distribution (sum of $k$ IID exponential random variables with mean $1$); $\tau_{k+1}-\tau_{k}$ is another exponential random variable with mean $1$. Therefore, $\ln\tau_{k+1}-\ln\tau_{k}\geq 0$ almost surely. What is more, the tail of the exponential distribution with mean 1 is given by:
(49) $\mathbb{P}(\tau_{k+1}-\tau_{k}\geq v)=e^{-v},\quad v\geq 0.$
Using (49), we get: for every $u\geq 0$,
(50) $\begin{split}\mathbb{P}&(\ln\tau_{k+1}-\ln\tau_{k}\geq u)=\mathbb{P}(\tau_{k+1}\geq e^{u}\tau_{k})\\ &=\mathbb{P}(\tau_{k+1}-\tau_{k}\geq(e^{u}-1)\tau_{k})=\mathbb{E}(\mathbb{P}(\tau_{k+1}-\tau_{k}\geq(e^{u}-1)\tau_{k}\mid\tau_{k})),\end{split}$
Next, the Laplace transform of the Gamma random variable $\tau_{k}$ with shape $k$ is
(51) $\mathbb{E}\left[\exp(-v\tau_{k})\right]=(1+v)^{-k}.$
Use independence of $\tau_{k+1}-\tau_{k}$ and $\tau_{k}$ for (50) and (51):
(52) $\mathbb{P}(\ln\tau_{k+1}-\ln\tau_{k}\geq u)=\mathbb{E}\left[\exp(-\tau_{k}(e^{u}-1))\right].$
Next, apply the Laplace transform of $\tau_{k}$ to $v:=e^{u}-1$. The right-hand side of (52) is $(1+(e^{u}-1))^{-k}=e^{-ku}$. This is the tail of the exponential distribution with mean $1/k$. ∎
(a) Upper Left End
(b) Lower Right End
Figure 7. Left panel: $(-\ln k,-\ln\tau_{k})$ for $k=1,\ldots,100$ and $y=-x$ for $x\in[0,\ln 100]$. Right panel: $(\ln(501-k),-\ln\tau_{k})$ for $k=1,\ldots,100$ and $y=\ln(500-e^{x})$ for $x\in[\ln 401,\ln 500]$. For each panel, we do three simulations (solid) and plot the deterministic function (dotted).
The next lemma is the key to our analysis of both ends of the standard normal curve (41). It is proved very similarly to our results in [27, Proposition A1], but we provide the full proof for completeness.
Lemma 4.
With probability $1$, the sequence $|\ln\tau_{k}-\ln k|,\,k=1,2,\ldots$ is bounded.
Proof.
From [2, Chapter 8, Problem 8.7], we know that the random variables $\ln\tau_{k}-\ln\tau_{k-1}$ are independent. By Lemma 3, the mean of $\ln\tau_{k+1}-\ln\tau_{k}$ is $1/k$ and the variance is $1/k^{2}$. These differences $\ln\tau_{k+1}-\ln\tau_{k}$ all are independent. Therefore,
(53) $\mathbb{E}\left[\ln\tau_{k}\right]=\sum_{j=1}^{k}\frac{1}{j}=:S(k),\quad\mathrm{Var}\left[\ln\tau_{k}\right]=\sum_{j=1}^{k}\frac{1}{j^{2}}:=V(k).$
It is well-known that $S(k)\sim\ln k$ as $k\to\infty$, and the sequence $(S(k)-\ln k)$ is bounded. Also, $V(k)\to\pi^{2}/6$ as $k\to\infty$. By [26, Theorem 1.4.2], from (53) we complete the proof. ∎
Thus we can replace $\ln\tau_{k}$ in (48) and (47) with $\ln k$. This result allows us to approximate curves (47) and (48) with continuous functions. The upper-left end of the curve (47) then becomes $(\ln k,-\ln k)$. This is a linear curve with 45 degrees of incline. Next, the lower-right end of the curve (48) becomes $(\ln k,\ln(N+1-k)$. Let $x=\ln k$ and $y=\ln(N+1-k)$: Then
$y=\ln(a-e^{x}),\quad a:=N+1.$
This function is concave but obviously not linear:
$y^{\prime\prime}=\left(\frac{(a-e^{x})^{\prime}}{a-e^{x}}\right)^{\prime}=\left(1+\frac{a}{e^{x}-a}\right)^{\prime}=-\frac{ae^{x}}{(e^{x}-a)^{2}}<0.$
This reproduces the shape of this capital distribution curve. See Figure 7 for the upper left and lower right ends of the capital distribution curve. The simulation of $\ln\tau_{k}$ for $k=1,\ldots,100$ and the plot of the comparative deterministic function (with $N=500$) was done in Python. The code standard-curve-simulation.py is in GitHub repository asarantsev/size-capm-vix The shape of both ends of the curve in Figure 6 are reproduced here in Figure 7.